# Global-Local Interface with Selective Direct and Singularity-Avoiding Motion Mapping for Intuitive Teleoperation

### v2

@@ V Operational Jacobian in Teleoperation
VI Experimental Validations
This section presents experimental results validating the performance of G-L interface leveraging the direct motion mapping (Section III) and singularity-avoiding motion mapping (Section IV). The evaluation focuses on three key aspects:
  1. 1.
    Precision Control: Using direct motion mapping, we assessed the system’s capability for high-accuracy telemanipulation through tasks with stringent precision requirements.
  2. 2.
    Workspace Coverage: Also employing direct motion mapping, we evaluated the ability to command extensive movements across the slave robot’s workspace, demonstrating effective utilization of its kinematic limits.
  3. 3.
    Orientation Dexterity and Singularity Handling: Utilizing the singularity-avoiding mapping, we demonstrated intuitive control during tasks involving significant orientation changes, specifically highlighting smooth trajectory execution during the traversal of wrist singularities.
In addition to these demonstrations, a quantitative comparison was conducted to assess the benefits of the proposed integrated Global-Local (G-L) interface. Task completion times for both fine-grained manipulation and large-range motion scenarios were measured using the full G-L interface and compared against performance when using only the global or local component in isolation. This analysis serves to quantify the advantages derived from combining these complementary components.
VI-A Implementation Details
The global components are actuated with Dynamixel XL330-M288 servomotors arranged to a desired joint-link chain mimicking the Pieper robot arms. For direct mapping, the global component has full 6 DoFs, and for local mapping, it has only 3 DoFs. Fig. 5 shows snapshots of G-L interface realization. The motion correspondence of the direct mapping and sigularity-avoiding realizations can be found in the supplementary video. The local component comprises two 3D System 6-DoF Touch devices for precise end-effector control in Cartesian space. The isolated global components in the comparison experiments has 6 DoFs.
VI-B Precise Manipulation Demo
Figure 7: Snapshots of tasks requiring large workspace. (A) Move an object from the desk to the cabinet. (B) Take a bag from the cabinet and place it on the desk. (C) Bimanual handling of oranges from the front-left of the workspace to the back-right. (D) Bimanual transferring a test tube from a wash basin to a rack. (E) Pouring the liquid from a test tube into a beaker, and then place the test tube back on the rack.
To assess the precision and dexterity afforded by the direct mapping teleoperation realization, we conducted a series of experiments involving intricate manipulation tasks. These tasks, detailed below, were selected to represent challenges commonly encountered in real-world scenarios and to highlight the system’s capabilities in addressing them. Snapshots of these tasks are shown in Fig. 6.
Figure 8: Snapshots of daily object bimanual manipulation. (A) Opening a snack box. (B) Opening bottle cap. (C) Opening chip bag.
@@ VI-B Precise Manipulation Demo
Figure 9: Comparison of completion time between global component only, local component only, and our proposed G-L interface. The G-L interface outperforms only global component in most tasks, except in the Nested Bucket Picking. As shown in Table I, despite taking less time to complete successful trials, only global component has a significantly lower success rate in Nested Bucket Picking. In large distance motion, G-L interface significantly outperforms local component only.
@@
VI-C Large Linear Range Manipulation Demo
TABLE I: Success rates of isolated global component and G-L interface across different tasks. Tasks are selected based on predefined failure criteria described in Section VI-B.
Task Global Component G-L Interface
Nested Bucket Picking $3/7$ $6/7$
Pencil Picking $/7$ 7 $/7$ 7
Membrane Touching $5/7$ $/7$ 7
Buckle Opening $6/7$ $/7$ 7
Handle Lifting $/7$ 7 $/7$ 7
VI-D Large Rotational Range Manipulation Demo
To assess the ability of the singularity avoidance mapping on intuitive orientation control and robustness against wrist singularities, we conducted demonstrations involving considerable end-effector orientation change and needs to cross the wrist singularities.
  • Test Tube Pick and Place: This demonstration involves picking up a flat-laying test tube from a basin, handle it over to another robot arm and place it on a rack. This task requires significant pitch angle change and passes through 156 and 456 singularities.
  • Liquid Pouring: This demonstration involves picking up a test tube from the rack, pour the liquid into a beaker and place the test tube rack. This task requires significant yaw angle change.
Both tasks are completed efficiently and their snapshots are shown in Fig. 7D and 7E. These tasks are good examples of how singularity avoidance motion mapping yields intuitive orientation control and robustness against wrist singularities.
VI-E Daily Object Manipulation
We additionally show the G-L interface can also complement bimanual teleoperation tasks involving daily objects as in Fig. 8. These tasks include picking up a plate, where one slave robot must push down one edge to lift up the other, thereby allow another slave robot to pick up the entire plate. Tasks like opening a chip bag, snack box, and water bottle are also included.
VI-F Fine-Grained Task Completion Time Comparisons
To quantitatively assess the performance improvements enabled by the G–L hierarchical teleoperation system, we conducted a comparative study involving three control conditions: (1) using the global component only, (2) using the local component only, and (3) using both the global and local components. Seven participants were recruited to perform a series of representative tasks, each completed twice under every condition. Regrasping and retrying were allowed unless the task was classified as a failure based on predefined criteria (detailed below). Failed trials were excluded from the computation of average completion time.
Prior to the experiment, all participants received a brief tutorial on system operation and were allotted one hour of practice to familiarize themselves with the interface and task procedures. For each trial, the master and slave systems were reset to identical initial configurations, and the task completion time was recorded.
@@ VI-F Fine-Grained Task Completion Time Comparisons
  • Handle Lifting: Participants rotate a laid-down box handle to an upright position and lift the box. Failure conditions are identical to the Buckle Opening task.
  • Dots Aiming: Abstracting common insertion tasks (e.g., SIM ejector, shafts), participants use the slave gripper to prick three target dots (1 mm and 3 mm radius) with a needle. The target sheet’s location is hidden during trials but available for practice. No failure criterion; continuous attempts are allowed until success.
  • Large Distance Motion: This task compares the G–L interface with a local-only teleoperation setup to evaluate efficiency in executing large-range movements. When repeatedly performing such motions, the G–L interface demonstrates a clear advantage in operational efficiency. Specifically, the G–L interface achieves a significantly shorter task completion time, averaging only $3.30\pm 0.53$ seconds, compared to $15.1\pm 1.3$ seconds using the local component alone.
@@
Fig. 9 presents the results of the quantitative study, demonstrating that the G–L interface consistently outperforms the isolated global or local components across most tasks. While participants using all configurations were able to quickly move the slave robot’s end-effector near the target area, those operating with the global-only setup frequently struggled to perform the final fine motions required for task completion. In contrast, the local component within the G–L interface enabled smoother and more efficient fine-tuning, resulting in significantly shorter overall completion times. The isolated global component system demonstrated marginally faster completion in nested bucket picking. However, it has a considerably lower success rate comparing to the G-L interface, as elaborated in Table I.
VII Conclusion, Limitations, and Future Work
This paper presents the Global–Local (G–L) teleoperation interface, which decouples large-scale positioning and fine manipulation through a hierarchical dual-mode control framework. We implemented both direct and singularity-avoiding motion mapping strategies, and introduced the concept of an operational Jacobian to analyze the smoothness and stability of joint behavior during local control. Experimental results across a range of tasks—including precision placement, large workspace traversal, and bimanual manipulation—demonstrate that the G–L interface improves efficiency, task success rate, and user experience compared to conventional global-only or local-only systems.
Nonetheless, several limitations remain. Manual mode switching introduces operational friction and may increase cognitive load. The current singularity-avoiding mapping strategy is task-driven and heuristic, lacking general optimality or theoretical uniqueness. Additionally, the system does not yet support force–torque feedback, and has not been validated under remote or delayed network conditions. Furthermore, the G–L interface’s modular architecture supports flexible combinations of global and local master devices, making it broadly adaptable to various application scenarios beyond the current implementation.
Future work will focus on integrating automatic or learning-based switching mechanisms, increasing the system’s degrees of freedom to improve singularity robustness, and incorporating visual and force–torque feedback for closed-loop, compliant control. We also plan to extend the operational Jacobian framework to accommodate non-Pieper structures and redundant manipulators, enabling broader applicability and more adaptive mapping strategies. These developments aim to further enhance the robustness, responsiveness, and scalability of the G–L teleoperation interface for real-world deployment in industrial, surgical, and exploratory environments.

### v3

@@ III Implementation Details
The single-arm teleoperation system was developed based on spatial decoupling. Its slave system comprises a Flexiv Rizon robotic arm equipped with an Inspire dexterous hand. The global component of the master device utilizes a Dynamixel servomotor structure to control the robot arm’s primary joints, while the local component employs two MTI-3 inertial measurement units (IMUs) and a dexterous exoskeleton for fine manipulation. The Inspire hand, a 6-DOF, five-fingered anthropomorphic hand, is mounted to the Flexiv arm’s end-effector with a 90-degree offset to mitigate kinematic singularities. The master system’s global component consists of four serially-linked Dynamixel XL330-M288 servomotors, kinematically replicating the Flexiv arm’s proximal four joints. For the local component, two MTI-3 IMUs are attached to the operator’s forearm and palm, respectively, and a dexterous exoskeleton, integrated with an operational glove, captures index and thumb motions. All devices are interfaced and coordinated through a multi-process LabVIEW program. Upon first launch, the starting orientation of the two IMUs will be respectively recorded as the home orientation, denoted as $R^{i}_{1}$ and $R^{i}_{2}$. When the local component is activated and the two IMUs report current orientations $R_{1}$ and $R_{2}$, we take their rotational displacements ${R^{i}_{1}}^{-1}R_{1}$ and ${R^{i}_{2}}^{-1}R_{2}$. The difference between the forearm and hand IMU, which is $R_{s}=R_{1}^{-1}R^{i}_{1}{R^{i}_{2}}^{-1}R_{2}$, represents the combined rotation of the last three joints in the Flexiv arm. The last three joint positions are computed by finding the instrinsic X-Y-Z Euler angles that lead to $R_{s}$.
An exoskeleton is designed for the single-arm teleoperation system to control the Inspire Hand. It is equipped with six AS5600 magnetic encoders and can be attached to the operator’s hand to record the thumb and index flexion of the human operator, as shown in Fig. 6. The thumb of Inspire Hand will copy the motion of the operator’s thumb, whereas the index, middle, ring, and pinky fingers will imitate the motion of the operator’s index following the virtual finger concept [37, 38]. The schematics and mapping methods of the exoskeleton will be detailed in the supplementary materials.
IV Experimental Validations
@@
Figure 8: Comparison of completion time between global component only, local component only, and our proposed G-L interface. The G-L interface outperforms only global component in most tasks, except in the Nested Bucket Picking. As shown in Table II, despite taking less time to complete successful trials, only global component has a significantly lower success rate in Nested Bucket Picking. In large distance motion, G-L interface significantly outperforms local component only.
The performance of the implemented temporal decoupling and spatial decoupling teleoperation systems, mentioned in section II, are evaluated through three distinct task categories. The first category comprises tasks with stringent accuracy requirements designed to assess the master system’s capacity for precise telemanipulation. The second category encompasses tasks necessitating extensive movement relative to the slave system’s workspace, demonstrating the system’s ability to effectively utilize the robot’s full range of motion. Finally, the third category involves tasks demanding the control of specialized dexterous end-effectors, specifically a five-fingered robotic hand in this study.
IV-A Precise Manipulation
Figure 9: Snapshots of tasks requiring large workspace. (A) Move an object from the desk to the cabinet. (B) Take a bag from the cabinet and place it on the desk. (C) Bimanual handling of oranges from the front-left of the workspace to the back-right. The supplementary video of this paper shows the whole process of these demonstrations.
To assess the precision and dexterity afforded by the temporal decoupling teleoperation system (dual-arm setup), we conducted a series of experiments involving intricate manipulation tasks. These tasks, detailed below, were selected to represent challenges commonly encountered in real-world scenarios and to highlight the system’s capabilities in addressing them. Snapshots of these tasks are shown in Fig. 7.
@@ IV-A Precise Manipulation
To quantitatively evaluate the performance improvements offered by the temporal decoupling teleoperation system, we conducted a comparative study with isolated global and local component. Seven participants were recruited to complete a series of tasks under two conditions: (1) using the global component (equivalent to the GELLO system) or local component only, and (2) using both the global and local components. Each participant performed each task twice, once under each condition. Regrasping and retrying were permitted unless the task was deemed failed according to the predefined criteria detailed below. In the event of failure, the respective trial was disregarded when computing the average completion time. Prior to the experiment, participants received a brief tutorial on operating the teleoperation system and were given one hour of practice time to familiarize themselves with the system and task procedures. For each trial, the master and slave systems were initialized to identical starting configurations, and the completion time was recorded.
@@
  • Handle Lifting: Given a box whose handle is initially laid down, participants are required to rotate the handle to an upright pose and lift the box using the handle. The failure condition is the same as the Buckle Opening task.
  • Dots Aiming: This task is an abstraction of many accurate insertion tasks like SIM card needles, shafts, and TRS connectors. Three dots with radius 1 mm, 2 mm, and 3 mm are printed on a target sheet. The slave system pinches a needle with its gripper. Volunteers are asked to prick each dot with this needle using the master system. Participants cannot access the test time target sheet position, but they are allowed to practice with the same target sheet before the test starts. This task is not given a failure criteria so participants can make continuous attempts until they are successful.
Figure 10: Snapshots of daily object manipulation. (A) Opening a snack box. (B) Dual arm plate grasping. (C) Opening bottle cap. (D) Opening chip bag.
@@
Fig. 8 presents the results of the quantitative study, demonstrating that the G–L interface consistently outperforms the isolated global or local components across most tasks. While participants using all configurations were able to quickly move the slave robot’s end-effector near the target area, those operating with the global-only setup frequently struggled to perform the final fine motions required for task completion. In contrast, the local component within the G–L interface enabled smoother and more efficient fine-tuning, resulting in significantly shorter overall completion times. The isolated global component system demonstrated marginally faster completion in nested bucket picking. However, it has a considerably lower success rate comparing to the G-L interface, as elaborated in Table II.
Global Component G-L Interface
Nested Bucket Picking $3/7$ $6/7$
Pencil Picking $7/7$ $7/7$
Membrane Touching $5/7$ $7/7$
Buckle Opening $6/7$ $7/7$
Handle Lifting $7/7$ $7/7$
TABLE II: Success rate of isolated global component and G-L interface on each task. Only tasks with pre-defined failure conditions in section IV-A are included. The G-L interface shows a higher success rate
IV-B Large Range Manipulation
Figure 11: Snapshots of tasks with dexterous manipulation. (A) The operator needs to fetch a box of drink from a distance away, take a sip and then put the drink box back. (B) Reorientating a water bottle from the desk.
IV-C Daily Object Manipulation
Tasks demonstrated in sections IV-A and IV-B are mainly achieved via single-arm manipulation. In this section, we show that the temporal decoupling teleoperation system can also complement bimanual teleoperation tasks involving daily objects. These tasks include picking up a plate, where one slave robot must push down one edge to lift up the other, thereby allow another slave robot to pick up the entire plate. Tasks like opening a chip bag, snack box, and water bottle are also included.
IV-D Dexterous Manipulation
This section details the master system built from spatial decoupling that is capable of controlling a complex end-effector. We use two tasks to demonstrate its dexterity: Feeding the human operator a drink and reorientating a water bottle. Snapshots of these tasks are shown in Fig. 11.
V Discussions and Limitations
Despite its advantages, the proposed G-L interface presents certain limitations, primarily stemming from its design and hardware implementation. These limitations, outlined below, warrant consideration for future development and refinement.
  • Additional Hardware Costs: Unlike conventional teleoperation systems that typically employ a single master device per slave robot, the G-L interface often requires two sets of master devices. While this configuration contributes to enhanced performance, it also leads to increased hardware costs.
  • Bimanual Control for Unimanual Tasks: The temporal decoupling method inherent to the G-L interface allows for straightforward bimanual teleoperation by preventing simultaneous activation of the global and local components. However, in the spatial decoupling method, where both components are engaged concurrently, the operator must utilize both hands to control a single slave robot, potentially increasing operational complexity.
  • Challenges in Orientation Control: Empirical observations revealed that fine-tuning the orientation of the slave robot’s end-effector proved more challenging than adjusting its position. Within the temporal decoupling framework, the local component’s reliance on scaled motion tracking from the touch-X device limits its efficacy to small orientation adjustments. Larger adjustments can become unintuitive due to the scaled mapping between master and slave rotations. Conversely, while the dual-IMU design in the spatial decoupling framework permits larger orientation adjustments, it introduces sensitivity to unintended skin movement as the operator twists their arm. These movements can increase discrepancies between the measured wrist rotation and its ground truth, leading to counterintuitive behavior in the slave end-effector’s orientation. This issue could potentially be mitigated by incorporating a more robust and precise wrist rotation sensing mechanism.
Interestingly, a notable learning curve was observed among participants. Initial trials were often characterized by cautious and slow movements, with proficiency and speed increasing with practice. To account for this learning effect and ensure a fair comparison, each participant was provided with one hour of dedicated practice time prior to data collection.
VI Conclusion
This paper introduced the G-L interface for on-demand teleoperation, a novel approach that decouples the master system into global and local components responsible for large-range movements and fine manipulations, respectively. This design enables the master system to effectively make precise adjustments while retaining the capacity for utilizing the entire workspace of the slave robot. The modular nature of the G-L interface allows for flexible implementation tailored to specific task requirements.
This work presented two realizations of the G-L interface: temporal decoupling and spatial decoupling. Through demonstrations and experimental evaluation, we have shown that temporal decoupling facilitates tasks demanding high accuracy while maintaining full workspace utilization. Conversely, spatial decoupling enables the teleoperation of complex end-effectors with enhanced dexterity. These findings highlight the versatility and potential of the G-L interface to advance the capabilities of teleoperation systems across a wide range of applications.
Appendix A: Exoskeleton Design
Figure 12: Design details of the exoskeleton master that controls the Inspire dexterous hand.
For better fit with the human operator’s hand, we added a $60^{\circ}$ tilting angle to the thumb of the exoskeleton. Fingers on the exoskeleton have two joints, each equipped with an AS5600 magnetic encoder. We designed a circuit to send encoder readings to a multiplexer via I2C protocol. The multiplexer is then connected to an ESP32 microprocessor, which allows communication with a computer. Fig. 12 further elaborates the exoskeleton’s structural design and the communication circuit.
Figure 13: Mapping between the exoskeleton and Inspire dexterous hand’s state. Subfigures (A), (B) and (C) are duplicated from the manual of the Inspire hand.
The Inspire dexterous hand is cable-driven with 6 DoFs in total, two on the thumb and one for each of the remaining fingers. Its position state is represented by the bending and rotational angles between the fingertip and the metacarpal plane. Therefore, while mapping the exoskeleton’s position state to the Inspire hand’s position state, we also compute the angle between the fingertip and metacarpal plane of the exoskeleton, as shown in Fig. 13.
Appendix B: Further Details of Tasks in Fig. 8
This part includes snapshots of tasks in Fig. 8 of the paper for better understanding of the testing criteria. Tasks that have already appeared in Fig. 7 and the dots aiming tasks are not included here.
(a) Nested Bucket Picking: Taking out the smaller bucket from the larger bucket. Reattempts commonly happens when both buckets are taken up, or the larger bucket is tipped over when lifting up the smaller bucket. Reattempts usually cause an uncontrollable motion of the larger bucket, and the task is considered failed if the large bucket fall off the platform.
(b) Pencil Picking: If the gripper hits the desk too hard, emergency stop will be triggered and the task is considered failed.
(c) Membrane Touching: The middle subfigure is a succeed case where the needle is touching the scotch tape. The right subfigure is a failure case where the needle penetrates through the scotch tape.
(d) Buckle Opening: If the gripper pushes the box too hard in a wrong direction, emergency stop will be triggered and the task is considered failed.
(e) Handle Lifting: the slave robot must rotate the and and lift up the box to be considered success.
Figure 14: Snapshots and further explanations of tasks for quantitative studies in Fig. 8.